OpenStudy (anonymous):

The length of a rectangle is four times its width. If the area of the rectangle is 100in^2 , find its perimeter.

OpenStudy (p0sitr0n):

L = 4w A = L*w = 4w*w = 4w^2=100 w^2=25 w=+-5, but distance cant be negative, so choose +5 P = 2L +2w = 2*4w + 2*w = 10w= 10*5 = 50 in